OpenStudy (anonymous):

evaluate -d4 when d = -8 simplify your ans. type an integer or a decimal.

OpenStudy (saifoo.khan):

32

OpenStudy (anonymous):

can u show me how to set this up for solving

OpenStudy (saifoo.khan):

sure

OpenStudy (anonymous):

\[ -d^4 = -(-8)^4 = -(4096) = -4096 \] or if you meant \(-d \cdot 4 \), then 32.